# Language Model Preference Evaluation with Multiple Weak Evaluators

### v2

@@ 3 Theoretical Analysis
In this section, we provide a theoretical foundation for our method, showing that by modeling preference graphs as random perturbations of a ground truth DAG, GED can reliably recover the true structure through graph ensemble and denoising with high probability, demonstrating its robustness in handling noisy evaluations.
Theoretically, we treat each of our preference graph as a random perturbation of some ground truth DAG $G=(V,A)$. Specifically, we consider a random graph generator $\mathcal{G}(G,\delta_{1},\delta_{2})$ with parameters $\delta_{1},\delta_{2}\in[0,1]$ such that $G_{i}=(V_{i},A_{i})\sim\mathcal{G}(G,\delta_{1},\delta_{2})$ satisfies $V_{i}=V$. Furthermore, for each $u,v\in V$ with $u\neq v$,
@@
    $\mathbb{P}((u\rightarrow v)\in A_{i})=1-\delta_{1}\quad\text{and}\quad\mathbb{P}((v\rightarrow u)\in A_{i})=\delta_{1};$
@@
    $\mathbb{P}((u\rightarrow v),(v\rightarrow u)\notin A_{i})=1-\delta_{2},\quad\mathbb{P}((u\rightarrow v)\in A_{i})=\frac{\delta_{2}}{2}\quad\text{and}\quad\mathbb{P}((v\rightarrow u)\in A_{i})=\frac{\delta_{2}}{2}.$
@@
$\mathbb{P}\left(G\subseteq\text{MAS}(\widehat{G})\right)\geq 1-2|A|\exp\left(-\frac{N\epsilon^{2}}{2}\right)-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right),$
@@
The full proof is given in Appendix E. From the theorem, we can see that the probability of failure decreases exponentially as the number of samples $N$ increases. Meanwhile, this guarantee only requires $\delta_{1}<0.5$ and does not place restrictions on $\delta_{2}$, which are very mild conditions.
4 Experiments on Response Selection
Table 1: Performance comparison of response selection methods across five benchmarks. GED consistently outperforms baseline methods, demonstrating the effectiveness of graph denoising and the aggregation of weaker evaluators.
@@ 4 Experiments on Response Selection
In this section, we evaluate the performance of GED on five benchmarks: HumanEval (Chen et al., 2021), AlpacaEval (Li et al., 2023b), MATH (Hendrycks et al., 2021), GSM8k (Chen et al., 2021), and GAIA (Mialon et al., 2023). The Qwen2-72B (Yang et al., 2024a) model ($\mathcal{M}$) generates ten candidate responses per question, and we assess the effectiveness of different methods in selecting the best response. For further implementation details, see Appendix A. We evaluate performance using three setups. First, in the single model setting, the baselines include ContraSolver(Zhang et al., 2024b), Self-consistency(Wang et al., 2022), and direct evaluation with models (Llama3-8B, Mistral-7B, Qwen2-7B and Qwen2-72B). Additionally, we include a baseline called ListPreference, where instead of pairwise comparisons, all candidate responses are input into Qwen2-72B for selecting the most appropriate response. Then, in the single evaluator setting, individual evaluators (Llama3-8B, Mistral-7B, Qwen2-7B, Qwen2-72B) select the best response from $\mathcal{M}$’s outputs, with and without applying GED’s graph denoising. Finally, in the multiple weak evaluators setup, we combine three weaker evaluators (Llama3-8B, Qwen2-7B, Mistral-7B) to select responses from Qwen2-72B with GED. We also include a new baseline referred to as Multi-MV, which just takes the majority for each pairwise comparison as the preferred answer. We present the results of GED and its variant (w/o denoising), which ensembles the preference graphs without the denoising step.
Main results.
Table 1 presents the results of the response selection task across five benchmarks. GED consistently outperforms baseline methods, including both single model evaluations (single model) and direct response selection by individual models (single evaluator). This demonstrates the strength of aggregating weak evaluators with GED, particularly when coupled with graph denoising, which enhances response quality by filtering out noise and biases. This highlights the effectiveness of aggregating weak evaluators and applying graph denoising to improve response quality. Furthermore, by combining preference graphs from weaker models (Llama3-8B, Mistral-7B, Qwen2-7B), GED surpasses the performance of a much stronger evaluator (Qwen2-72B). This underscores the value of ensemble methods in mitigating the limitations of individual evaluators. Then, the denoising process proves to be crucial for improving consistency and overall response quality. The substantial performance gains observed when using GED with denoising, compared to both the single evaluator setup and the ensemble without denoising, highlight its importance in refining response selection. We also evaluated Multi-MV, which aggregates pairwise comparisons using majority voting across evaluators. While Multi-MV offers improvements over individual evaluators, it underperforms compared to GED, highlighting GED’s ability to capture more nuanced evaluation signals and reduce inconsistencies. Additionally, we observed that the ListPreference baseline performed worse than Qwen2-72B as single evaluator, likely due to LLM limitations in handling long-text. Lastly, to further evaluate GED, we compared its performance with GPT-3.5, GPT-4-o-mini, and GPT-4-o. Due to computational and API cost constraints, we limited the evaluation to 100 data points for each task. As shown in Figure 3, GED consistently outperformed GPT-3.5 across all tasks and surpassed GPT-4-o-mini on challenging benchmarks like HumanEval and GSM8k. These results highlight the superiority of GED, particularly in leveraging multi-weak evaluators and graph denoising to outperform individual state-of-the-art models.
Ablation study.
We evaluate the impact of removing the ensembling step in GED, referred to as the (w/o ensemble) variant. In this case, individual evaluators’ preference graphs are denoised and converted to rankings, which are then aggregated using methods such as Weight Score, Kemeny, Weighted Kemeny, Pairwise Majority, and Weighted Pairwise Majority (detailed in Appendix M). For simplicity of presentation, we use Weight Score to represent GED (w/o ensemble) (Weight Score). As shown in Figure 4, all (w/o ensemble) methods consistently underperform compared to GED. This performance gap arises because converting graphs to ranks before aggregation leads to information loss. In contrast, GED ensembles the graphs directly, preserving more detailed preference information and resulting in better final rankings.
@@ 5 Experiments on Model Ranking
In this section, we evaluate the effectiveness of GED in the model ranking task within a human preference setting, using the AlpacaEval benchmark (Li et al., 2023b). We employ 30 widely used models from the AlpacaEval dataset as our model set $\mathcal{M}$, while the benchmark’s questions form the question set $Q$. The rankings provided by the AlpacaEval benchmark serve as ground truth for evaluating the accuracy of various ranking methods. This is justified by AlpacaEval’s strong correlation with Chatbot Arena rankings, making it a reasonable proxy for human judgments (Dubois et al., 2024a). We adopt Ranking Correction, measured by the Spearman rank correlation coefficient, to evaluate the similarity. To generate rankings, we utilize outputs from the open-source models Llama3-70B, Qwen2-72B, Mistral-8$\times$7B, and Qwen1.5-72B as our evaluators, denoted as set $\mathcal{A}$. For further implementation details, see Appendix A. We investigate two variants of GED: (w/o ensemble) denoises the preference graphs from different evaluators for the same question, converts each into a ranking, and then ensembles these rankings to produce the final output, while (w/o denoising) directly ensembles the preference graphs to obtain the final ranking without denoising.
@@
6 Experiments on Instruct Tuning
Experiment Setup.
In this section, we explore the effects of various data selection methods for model alignment on Llama-2-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023) through instruct tuning. Specifically, we randomly sampled 5000 data points from UltraFeedback (Cui et al., 2023), and utilized Qwen1.5-14B (Yang et al., 2024a) to generate 8 responses for each data point as instruct data. We then applied four different methods—Random, Longest (Zhao et al., 2024), ContraSolver (Zhang et al., 2024b), and our proposed GED—to select a subset of these responses for model alignment training. The Origin refers to the performance of the base model without alignment. The resulting models were evaluated on the HH-RLHF (Bai et al., 2022) benchmark, which consists of four sub-sets: Harmless (base), Helpful (base), Helpful (online), and Helpful (rejection). For evaluation, we adopted the same Reward model, following prior work (Song et al., 2024; Yu et al., 2023), to measure human preference levels gained by the models. These results are summarized in Table 3. To ensure a comprehensive evaluation, we also tested the models, using the Llama-2-7B backbone, across additional benchmarks such as LIMA (Zhou et al., 2023), Vicuna (Chiang et al., 2023), Koala (Vu et al., 2023), WizardLM (Xu et al., 2023) and Self-Instruct (Wang et al., 2022), in line with recent works (Chen et al., 2023b; Zhang et al., 2024a; Hu et al., 2024).
@@ 6 Experiments on Instruct Tuning
From Table 3, we observe that GED consistently outperforms all baseline methods, demonstrating its effectiveness in selecting high-quality responses when multiple answers are available for a given instruction. When faced with multiple responses $y_{1},y_{2},\ldots,y_{n}$ for a given instruction $x$, the Random selection method can have a detrimental impact, especially when the quality of the responses is inconsistent. This effect is most evident with the Mistral-7B, where Random selection actually performs worse than the Origin, indicating that randomly chosen data points can introduce noise and degrade the model’s performance. Moreover, we find that simply selecting the longest response does not always lead to the best outcomes. While longer responses may provide more detailed answers, they are not necessarily better in terms of quality, particularly when both high-quality and low-quality answers exist for the same question. This is reflected in the results where the Longest method underperforms compared to both ContraSolver and GED, emphasizing that response length alone is not always a reliable criterion. From Figure 5, we can draw similar conclusions as Table 3. Specifically, we observe that GED consistently outperforms all baselines, demonstrating its effectiveness across all datasets. Particularly in AlpacaEval and Self-Instruct, the Random baseline performs worse than the Origin model, indicating that when response quality varies significantly, poor selection can lead to negative performance. In contrast, GED excels by aggregating preference graphs and applying denoising, effectively filtering out low-quality responses. This ensures robust performance, especially in cases where response quality is inconsistent. The denoising step is crucial for removing noisy evaluations, leading to improved model alignment. The denoising process in GED proves essential, particularly in settings with inconsistent responses, as it removes evaluation noise and leads to more robust performance.
7 Related Work
Preference evaluation of LLMs.
Reference-free evaluation metrics have a long history (Louis & Nenkova, 2013), which evaluates the generated text based on intrinsic properties and coherence with the context. Although they achieve high accuracy on matching inner-evaluator, the achievement suffers from spurious correlations such as perplexity and length (Durmus et al., 2022). Recently, people have started using a strong model (e.g., GPT-4) as an evaluator to perform a zero-shot reference-free evaluation on the weak models (Shen et al., 2023; Dubois et al., 2024b; Chen et al., 2023b). However, using LLM-based preference evaluations can introduce inconsistencies in preference graphs, often resulting in cyclic preferences or contradictions when comparing multiple outputs.
Weak supervision.
The concept of weak-to-strong supervision originates from the need to leverage noisy or partial labels in machine learning tasks, enabling the development of more robust models from imperfect data (Ratner et al., 2016; Zhang et al., 2023b; 2022). In LLMs, weak-to-strong supervision aids AI alignment by allowing weaker models to improve strong ones, enhancing performance without extensive data and supporting scalable oversight (Zheng et al., 2024a; Guo & Yang, 2024; Tong et al., 2024). Similarly, in task-oriented LLMs, weak-to-strong learning improves LLM’s ability by enabling strong models to refine their data autonomously, boosting performance without extensive high-quality input (Zhang et al., 2023a; Yang et al., 2024b). Through weak-to-strong supervision, LLM performance can be significantly improved by iteratively transforming low-quality labels into more reliable ones, leading to more effective model training and robust outputs (Zakershahrak & Ghodratnama, 2024; Lang et al., 2024).
@@ 8 Conclusion
In this paper, we presented GED, a framework designed to address inconsistencies in pairwise preference evaluations by LLMs. By employing graph ensemble techniques and denoising, GED reduces cyclic patterns and enhances the reliability of evaluation outcomes. Our theoretical analysis shows that GED can recover the ground truth DAG under reasonable conditions, improving consistency in preference rankings. Extensive experiments across response selection, model ranking, and instruct tuning demonstrate the efficacy of our method. GED consistently outperformed baseline methods in both single-evaluator and multi-evaluator settings, particularly in scenarios where combining weak evaluators led to superior results over stronger individual evaluators. Future work will explore extending GED to broader evaluation frameworks and applying its principles to more complex decision-making tasks, including multi-agent systems and human-AI interaction.
Acknowledgement
This project was partially supported by an OpenAI “Superalignment Fast” grant.
@@ A.1 Experimental Setup
All experimental procedures were conducted on a machine equipped with an AMD EPYC 7543 32-Core Processor, 512GB memory, 128 CPUs, and four 80GB NVIDIA A800 GPUs. The references to Llama-2-7B, Llama3-70B, Llama3-8B, Mistral-7B, Mistral-8×7B, Qwen2-7B, and Qwen2-72B in the main text refer to the specific models: Llama-2-7b-chat-hf, Meta-Llama-3-70B-Instruct, Meta-Llama-3-8B-Instruct, Mixtral-7B-Instruct-v0.3, Mixtral-8x7B-Instruct-v0.1, Qwen2-7B-Instruct, and Qwen2-72B-Instruct. We utilized the reward model oasst-rm-2-pythia-6.9b-epoch-1 following prior works (Song et al., 2024; Yu et al., 2023). Each experiment was repeated three times, and the average performance was reported as the final result. Our training script was adapted from the example provided in LlamaFactory (Zheng et al., 2024b), specifically https://github.com/hiyouga/LLaMA-Factory/blob/main/examples/train_full/llama3_full_sft_ds3.yaml. The training was configured with a batch size of 1 per device, gradient accumulation steps of 4, a learning rate of 1e-5, and he model was trained for 3 epochs, with warmup over 20 steps and a cosine learning rate scheduler. For generating diverse responses from LLMs, we followed the configuration in Yuan et al. (2024), setting $T=0.7$ and $p=0.9$. For tasks such as AlpacaEval (Dubois et al., 2024b), we used GPT-4-o unless stated otherwise.

### v3

@@ 4 Theoretical Analysis
In this section, we provide a theoretical foundation for our method, showing that by modeling preference graphs as random perturbations of a ground truth DAG, GED can reliably recover the true structure through graph ensemble and denoising with high probability, demonstrating its robustness in handling noisy evaluations. Theoretically, we treat each of our preference graph as a random perturbation of some ground truth DAG $G=(V,A)$. Specifically, we consider a random graph generator $\mathcal{G}(G,\delta_{1},\delta_{2})$ with parameters $\delta_{1},\delta_{2}\in[0,1]$ such that $G_{i}=(V_{i},A_{i})\sim\mathcal{G}(G,\delta_{1},\delta_{2})$ satisfies $V_{i}=V$.
Furthermore, for each $u,v\in V$ with $u\neq v$,
@@
    $\mathbb{P}((u\rightarrow v)\in A_{i})$ $=1-\delta_{1},$
    $\mathbb{P}((v\rightarrow u)\in A_{i})$ $=\delta_{1}.$
@@
    $\mathbb{P}((u\rightarrow v),(v\rightarrow u)\notin A_{i})$ $=1-\delta_{2},$
    $\mathbb{P}((u\rightarrow v)\in A_{i})$ $=\frac{\delta_{2}}{2},$
    $\mathbb{P}((v\rightarrow u)\in A_{i})$ $=\frac{\delta_{2}}{2}.$
@@
$\begin{split}\mathbb{P}\left(G\subseteq\text{MAS}(\widehat{G})\right)&\geq 1-2|A|\exp\left(-\frac{N\epsilon^{2}}{2}\right)\\ &\quad-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right),\end{split}$
@@
The full proof is given in Appendix G. From the theorem, we can see that the probability of failure decreases exponentially as the number of samples $N$ increases. Meanwhile, this guarantee only requires $\delta_{1}<0.5$ and does not place restrictions on $\delta_{2}$, which are very mild conditions.
Table 1: Performance comparison of response selection methods across five benchmarks. GED consistently outperforms baseline methods, demonstrating the effectiveness of graph denoising and the aggregation of multiple evaluators.
@@ 5 Experiments on Response Selection
In this section, we evaluate the performance of GED on five benchmarks: HumanEval (Chen et al., 2021), AlpacaEval (Li et al., 2023b), MATH (Hendrycks et al., 2021), GSM8k (Chen et al., 2021), and GAIA (Mialon et al., 2023). The Qwen2-72B (Yang et al., 2024a) model ($\mathcal{M}$) generates ten candidate responses per question, and we assess the effectiveness of different methods in selecting the best response. For further implementation details, see Appendix A. We evaluate performance using three setups. First, in the single model setting, the baselines include ContraSolver(Zhang et al., 2024b), Self-consistency(Wang et al., 2022), and direct evaluation with models (Llama3-8B, Mistral-7B, Qwen2-7B and Qwen2-72B). Additionally, we include a baseline called ListPreference, where instead of pairwise comparisons, all candidate responses are input into Qwen2-72B for selecting the most appropriate response. Then, in the single evaluator setting, individual evaluators (Llama3-8B, Mistral-7B, Qwen2-7B, Qwen2-72B) select the best response from $\mathcal{M}$’s outputs, with and without applying GED’s graph denoising. Finally, in the multiple evaluators setup, we combine three small evaluators (Llama3-8B, Qwen2-7B, Mistral-7B) to select responses from Qwen2-72B with GED. We also introduce a baseline, Multi-MV, which selects the response that receives the most votes from evaluators in pairwise comparisons. We present the results of GED and its variant (w/o denoising), which ensembles the preference graphs without the denoising step.
@@
Main results.
Table 1 presents the results of the response selection task across five benchmarks. GED consistently outperforms baseline methods, including both single model evaluations (single model) and direct response selection by individual models (single evaluator). This demonstrates the strength of aggregating weak evaluators with GED, particularly when coupled with graph denoising, which enhances response quality by filtering out noise and biases. Furthermore, by combining preference graphs derived from smaller models (Llama3-8B, Mistral-7B, Qwen2-7B), GED outperforms a much larger evaluator (Qwen2-72B). This underscores the value of ensemble methods in mitigating the limitations of individual evaluators.
Then, the denoising process proves to be crucial for improving consistency and overall response quality. The substantial performance gains observed when using GED with denoising, compared to both the single evaluator setup and the ensemble without denoising, highlight its importance in refining response selection. For Multi-MV, while it improves upon individual evaluators, it still underperforms GED, highlighting GED’s ability to capture nuanced evaluation signals and reduce inconsistencies. Additionally, we observed that the ListPreference baseline performed worse than Qwen2-72B as single evaluator, likely due to LLM limitations in handling long-text. Lastly, to further evaluate GED, we compared its performance with GPT-3.5, GPT-4-o-mini, and GPT-4-o. Due to computational and API cost constraints, we limited the evaluation to 100 data points for each task. As shown in Figure 2, GED consistently outperformed GPT-3.5 across all tasks and surpassed GPT-4-o-mini on challenging benchmarks like HumanEval and GSM8k. These results highlight the superiority of GED, particularly in leveraging multi-weak evaluators and graph denoising to outperform individual state-of-the-art models.
Ablation study.
We evaluate the impact of removing the ensembling step in GED, referred to as the (w/o ensemble) variant. In this case, individual evaluators’ preference graphs are denoised and converted to rankings, which are then aggregated using methods such as Weight Score, Kemeny, Weighted Kemeny, Pairwise Majority, and Weighted Pairwise Majority (detailed in Appendix N). For simplicity of presentation, we use Weight Score to represent GED (w/o ensemble) (Weight Score). As shown in Figure 3, all (w/o ensemble) methods consistently underperform compared to GED. This performance gap arises because converting graphs to ranks before aggregation leads to information loss. In contrast, GED ensembles the graphs directly, preserving more detailed preference information and resulting in better final rankings.
@@ 6 Experiments on Model Ranking
In this section, we evaluate the effectiveness of GED in the model ranking task within a human preference setting, using the AlpacaEval benchmark (Li et al., 2023b). We employ 30 widely used models from the AlpacaEval dataset as our model set $\mathcal{M}$, while the benchmark’s questions form the question set $Q$. The rankings provided by the AlpacaEval benchmark serve as ground truth for evaluating the accuracy of various ranking methods. This is justified by AlpacaEval’s strong correlation with Chatbot Arena rankings, making it a reasonable proxy for human judgments (Dubois et al., 2024a). We adopt Ranking Correction, measured by the Spearman rank correlation coefficient, to evaluate the similarity. To generate rankings, we utilize outputs from the open-source models Llama3-70B, Qwen2-72B, Mistral-8$\times$7B, and Qwen1.5-72B as our evaluators. For further implementation details, see Appendix A. We investigate two variants of GED: (w/o ensemble) denoises the preference graphs from different evaluators for the same question, converts each into a ranking, and then ensembles these rankings to produce the final output, while (w/o denoising) directly ensembles the preference graphs to obtain the final ranking without denoising.
@@
7 Experiments on Instruct Tuning
Experiment Setup.
In this section, we explore the effects of various data selection methods for model alignment on Llama-2-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023) through instruct tuning. Specifically, we randomly sampled 5,000 data points from UltraFeedback (Cui et al., 2023) and used Qwen1.5-14B (Yang et al., 2024a) to generate eight responses per data point as instruct data. We then applied four different methods—Random, Longest (Zhao et al., 2024), ContraSolver (using Qwen2-72B as the evaluator) (Zhang et al., 2024b), and our proposed GED, which leverages Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators—to select a subset of these responses for model alignment training. The Origin refers to the performance of the base model without alignment. The models were evaluated on the HH-RLHF (Bai et al., 2022) benchmark, which comprises four subsets: Harmless (base), Helpful (base), Helpful (online), and Helpful (rejection). For evaluation, we employed the same Reward model as in prior work (Song et al., 2024; Yu et al., 2023) to quantify human preference levels. The results are presented in Table 3. For further implementation details, see Appendix A. To ensure a comprehensive assessment, we further evaluated the models—using the Llama-2-7B backbone—on additional benchmarks, including LIMA (Zhou et al., 2023), Vicuna (Chiang et al., 2023), Koala (Vu et al., 2023), WizardLM (Xu et al., 2023), and Self-Instruct (Wang et al., 2022), in accordance with recent studies (Chen et al., 2023b; Zhang et al., 2024a; Hu et al., 2024). The corresponding results are summarized in Figure 4.
@@ 7 Experiments on Instruct Tuning
From Table 3, we observe that GED consistently outperforms all baseline methods, demonstrating its effectiveness in selecting high-quality responses when multiple answers are available for a given instruction. When faced with multiple responses $y_{1},y_{2},\ldots,y_{n}$ for a given instruction $x$, the Random selection method can have a detrimental impact, especially when the quality of the responses is inconsistent. This effect is most evident with the Mistral-7B, where Random selection actually performs worse than the Origin, indicating that randomly chosen data points can introduce noise and degrade the model’s performance. Moreover, we find that simply selecting the longest response does not always lead to the best outcomes. While longer responses may provide more detailed answers, they are not necessarily better in terms of quality, particularly when both high-quality and low-quality answers exist for the same question. This is reflected in the results where the Longest method underperforms compared to both ContraSolver and GED, emphasizing that response length alone is not always a reliable criterion. From Figure 4, we observe similar trends as in Table 3. GED consistently outperforms all baselines across various datasets, demonstrating its effectiveness in selecting high-quality responses. Notably, in AlpacaEval and Self-Instruct, the Random baseline performs worse than the Origin model, highlighting that when response quality varies significantly, poor selection can degrade model performance. In contrast, GED leverages preference graphs and denoising techniques to filter out low-quality responses, ensuring more robust and reliable performance, particularly in settings with inconsistent responses, as it removes evaluation noise and leads to more robust performance.
Table 4: Performance comparison on nuanced quality metrics (%). GED outperforms individual evaluators and random selection across Factuality, Relevance, Coherence, Inform., Helpful. and Validity metrics.
Random 86.73 87.91 92.47 77.62 17.48 48.92 68.52
Llama3-8B 88.59 89.91 94.41 79.77 18.48 50.92 70.35
Mistral-7B 89.10 90.29 94.85 79.95 18.55 51.13 70.65
Qwen2-7B 89.25 90.44 95.03 80.09 18.58 51.21 70.77
GED 94.73 95.91 97.36 86.62 19.48 55.92 75.00
8 Evaluating GED on More Metrics
To assess GED on more metrics, we evaluated its impact on nuanced quality aspects of LLM outputs, including factuality, relevance, coherence, informativeness, helpfulness, and validity. Since LLM tasks often require subtle judgments, GED’s adaptability is crucial. Following (Chen et al., 2023a), we used Llama3-70B to generate ten candidate responses per query and applied GED for response selection, with Llama3-8B, Mistral-7B, and Qwen2-7B as evaluators. Table 4 shows that GED consistently outperformed individual evaluators and random selection across all metrics. It improved factuality by  5 % points over the best evaluator, while also enhancing relevance, coherence, informativeness, and helpfulness. These results highlight GED’s ability to aggregate preferences effectively, capturing subtle qualities of generated content and improving overall reliability and utility in LLM assessments.
@@ 9 Conclusion
In this paper, we presented GED, a framework designed to address inconsistencies in pairwise preference evaluations by LLMs. By employing graph ensemble techniques and denoising, GED reduces cyclic patterns and enhances the reliability of evaluation outcomes. Our theoretical analysis shows that GED can recover the ground truth DAG under reasonable conditions, improving consistency in preference rankings. Extensive experiments across response selection, model ranking, and instruct tuning demonstrate the efficacy of our method. GED consistently outperformed baseline methods in both single-evaluator and multi-evaluator settings, particularly in scenarios where combining small evaluators led to superior results over larger individual evaluators. Future work will explore extending GED to broader evaluation frameworks and applying its principles to more complex decision-making tasks, including multi-agent systems and human-AI interaction.
Impact Statements
This work enhances the reliability of model-based preference evaluation, reducing inconsistencies in AI judgments. It improves fairness and transparency in AI evaluation, benefiting NLP applications and ethical decision-making frameworks.
@@ A.1 Experimental Setup
All experimental procedures were conducted on a machine equipped with an AMD EPYC 7543 32-Core Processor, 512GB memory, 128 CPUs, and four 80GB NVIDIA A800 GPUs. The references to Llama-2-7B, Llama3-70B, Llama3-8B, Mistral-7B, Mistral-8×7B, Qwen2-7B, and Qwen2-72B in the main text refer to the specific models: Llama-2-7b-chat-hf, Meta-Llama-3-70B-Instruct, Meta-Llama-3-8B-Instruct, Mixtral-7B-Instruct-v0.3, Mixtral-8x7B-Instruct-v0.1, Qwen2-7B-Instruct, and Qwen2-72B-Instruct. We utilized the reward model oasst-rm-2-pythia-6.9b-epoch-1 following prior works (Song et al., 2024; Yu et al., 2023). Each experiment was repeated three times, and the average performance was reported as the final result. Our training script was adapted from the example provided in LlamaFactory (Zheng et al., 2024b)  [^1] https://github.com/hiyouga/LLaMA-Factory/blob/main/examples/train_full/llama3_full_sft_ds3.yaml. The training was configured with a batch size of 1 per device, gradient accumulation steps of 4, a learning rate of 1e-5, and he model was trained for 3 epochs, with warmup over 20 steps and a cosine learning rate scheduler. For generating diverse responses from LLMs, we followed the configuration in Yuan et al. (2024), setting $T=0.7$ and $p=0.9$. For tasks such as AlpacaEval (Dubois et al., 2024b), we used GPT-4-o unless stated otherwise.